OpenStudy (anonymous):

simplify completely 3x^2 + 21x - 54/ x^2 + 3x - 54

OpenStudy (anonymous):

you have to begin by factoring the numerator and denominator

OpenStudy (anonymous):

3x^2 + 21x - 54=3(x-2)(x+9) x^2 + 3x - 54=(x-6)(x+9)

OpenStudy (anonymous):